The length of the foundation of a rectangular house is to be 5 meters more than its width. Find the length of the house if the perimeter must be 70 meters.
Answer by richard1234(7193)

Suppose the width is W and the length is W+5. Then,

2(W + (W+5)) = 70

W + (W+5) = 35

2W + 5 = 35 --> W = 15, L = W+5 = 20 meters